Suppose that the long-run price elasticity of demand for gasoline is -0.45. Assume that the price of gasoline is currently $4.00 per gallon, the quantity of gasoline is 140 billion gallons per year, and the federal government decides to increase the excise tax on gasoline by $1.00 per gallon. Suppose that in the long run the price of gasoline increases by $0.60 per gallon after the S1.00 excise tax is imposed. a. Using the midpoint formula, after the tax is imposed, the new quantity of gasoline demanded isillion gallons per year.(Enter your rosponse rounded to two decimal places.)
Answer #1

Ans: 130.55 billion gallon

Explanation:
Here, Ed = -0.45        [the absolute value is 0.45]

           original price = $4

            New price = $4.60

          % increase in price = (0.60 / 4) * 100 = 15%

So, % decrease in quantity demanded = 0.45 * 15% = 6.75%

Original quantity demanded = 140 billion gallon

New quantity demanded = 140 - (0.0675 * 140) = 130.55 billion gallon